OpenStudy (anonymous):

Find all solutions to the equation. cos^2x + 2 (cos x) + 1 = 0

OpenStudy (anonymous):

\[\cos^2x+ 2(cosx)+1=0\]

OpenStudy (anonymous):

@campbell_st mind helping me with this one? It's basically the same question as my last one but not solving by the interval of \[2,\pi\]

OpenStudy (campbell_st):

ok... so you are asked to find the general solution if you have \[x = \cos^{-1}(-1)\] and in the previous solution and as you habe shown here \[x = \pi\] the general solution for an undefined domain is \[x = \pm \cos^{-1}(-1) + 2n \pi\] where n is an integer because of the periodic nature of the cos curve, the solution will repeat every 2pi hope it makes sense